the time allowed for completing a certain piece of work is 40 days. By extending the time by 10 days we find that 25 fewer man will be needed. How many men are required to finish the work in 40 days?
Answer by edjones(8007)

Let x=number of men needed to finish job in 40 days
40 * x = 40x man-days required
50 * x-25 = 50(x-25) man-days
40x=50(x-25)
40x=50x-1250
-10x=-1250
x=125 number of men needed to finish job in 40 days.
.
Ed